# ExMAG: Learning of Maximally Ancestral Graphs

### v1

@@ Abstract
As one transitions from statistical to causal learning, one is seeking the most appropriate causal model. Dynamic Bayesian networks are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by its vertices, and weighted oriented edges suggest the strength of the causal relationships. When there are confounders, one would like to utilize both oriented edges (when the direction of causality is clear) and edges that are not oriented (when there is a confounder), yielding mixed graphs. A little-studied extension of acyclicity to this mixed-graph setting is known as maximally ancestral graphs.
@@ 1 Introduction
Simpson’s paradox is not a paradox, but empirically analysing causality without considering confounding factors in statistical analysis[28]. An interesting example is the Berkeley graduate admission paradox. The data show that women face higher difficulties when applying to graduate schools. Nevertheless, the reason is that women tend to apply to popular departments. In this example, academic ability can be a potential confounder since ability could influence the choice of department and the probability of admission. Controversially, [26] analyzed gene expression data from both human and mouse samples and concluded that data cluster more by tissue rather than by species. Within one year of the publication, [17] quickly realized that the previous paper integrated both previously published and newly collected human and mouse gene expression data without eliminating batch effects, which significantly caused the wrong results. The batch effect, stemming from different labs and sample processing protocols, was a key confounder that distorted the apparent species-based clustering of gene expression. Addressing this confounding variable through correction (using ComBat) revealed that tissue, not species, was the primary driver of gene expression patterns, emphasizing the importance of recognizing confounders in causal analysis. Therefore, recognizing and controlling for these confounders is crucial to avoid drawing incorrect conclusions about causality.
@@
Meanwhile, confounders represent a significant challenge in the process of causal inference, as they can cause catastrophe in discovering the relationships between observed variables across various fields. In the fields of financial econometrics, hidden states, such as regional heterogeneity, economic policy, or unobserved individual characteristics, can be addressed by using Difference-in-Differences (DiD) [15], Instrumental variables (IV) and Mendelian randomization (MR) to avoid incorrect conclusions about risk and returns [23]. Significantly, in biomedical science, confounders such as socio-economic status, age, or lifestyle factors can distort the true causal relationship between treatments and outcomes. Techniques such as propensity score matching [35] and instrumental variable approaches have been widely used to mitigate the effects of confounding in clinical trials and observational studies. In bioinformatics, particularly in genome-wide association studies (GWAS), confounders, including population stratification and environmental exposures, must be controlled to avoid biased estimates in genetic association studies [8, 18]. To mitigate such biases, statistical models that explicitly account for hidden confounders, such as spectral methods and latent variable models, are often employed [18]. Furthermore, meta-analysis and sensitivity analysis are often employed to evaluate the robustness of findings in the presence of potential confounders, especially when combining results from multiple studies [5, 27]. These methodologies ensure that the conclusions drawn are reliable and actionable, improving the credibility of statistical models across disciplines.
@@
In previous research, the effects of hidden bias in high-dimensional settings have been neglected. In [7] were confounders first given significant attention. Subsequently, this method leverages the dominant eigenstructure of time series to separate confounding influences and demonstrates that the performance of the estimation may improve as the dimension of the time series increases [25]. Recently, more statistical research has been introduced to enhance causal inference. Anchor regression, for instance, bridges the gap between causality and robustness by addressing heterogeneity in data [36]. Other significant contributions include spectral deconfounding, a technique designed to mitigate the effects of hidden confounders in high-dimensional settings [6]. This approach provides a framework for robust predictions in the presence of shifts in data distributions. Similarly, the invariance principle has emerged as a cornerstone of causal inference, linking causal structure to robust statistical models [5]. Furthermore, the concept of doubly robust inference offers an alternative framework for addressing hidden confounding factors, combining model robustness with efficiency in high-dimensional scenarios [18]. Together, these developments represent a significant leap forward in understanding and addressing the challenges posed by complex causal systems.
While previous work has focused primarily on constructing optimal Directed Acyclic Graphs (DAGs), these models have limitations when it comes to incorporating cycles and latent variables. On the other hand, Maximal Ancestral Graphs (MAGs) allow for more flexibility by accommodating both direct and indirect relationships among variables, even in the presence of confounding factors. In particular, MAGs can represent feedback loops and bidirectional relationships that DAGs cannot, making them a more powerful tool for capturing the complex dynamics of real-world causal systems. There are two classic types of graphical learning algorithms for constructing causal models: constraint-based and score-based methods, and sometimes hybrid methods that combine the first two approaches is considered as the third type. The canonical constraint-based method for learning DAGs (directed acyclic graphs) and DAGs (directed acyclic graphs) graphs are the PC and FCI algorithm, separately [39]. This type of method works magically in a way that conditional independences can be tested by statistical methods, and the graph can be reconstructed accordingly. However, a key limitation of constraint-based methods is that when the group of variables is large, errors in conditional independence tests tend to occur more often. Such errors will be caused in graph construction and mislead to a wrong causal relationship representing the data [31, 14].
Existing approaches to scoring ADMGs often rely on numerical methods for MLE estimation, as closed-form solutions are unavailable [32]. Additionally, factorization in ADMGs is not directly decomposable into individual variables and their parent sets, as in DAGs, but must instead consider components connected by bidirected paths (termed districts or c-components). The significant computational challenge of score-based methods is time-consuming, that is due to the number of graphs visited. In this way, [9] proposed that the Markov equivalence class (MEC) can be deployed to minimize scoring graphs in the visited MEC again. Our approach emphasizes the integration of MAGs with the recognition of confounders, offering a more accurate representation of causal relationships, as well as a more efficient method. By identifying the optimal MAG through Mixed-Integer Nonlinear Programming (MINLP), we can ensure that confounding factors are properly accounted for and that the resulting causal model reflects the true underlying data-generating process.
@@ 1.1 Motivating Example
Berkeley graduate admission paradox is famous in statistical teaching to illustrate Simpson’s paradox. Like most paradoxes, there is no violation of logic, just of intuition. And since different people have different intuitions, Simpson’s paradox means different things to different people. The poor intuition being violated in this case is that a positive association in the entire population should also be held within each department. Overall, females in these data did have a harder time getting admitted to graduate school. But that arose because females applied to the hardest departments for anyone, male or female, to gain admission to.
Perhaps a little more paradoxical is that this phenomenon can repeat itself indefinitely within a sample. Therefore, we proved that our method ExMAG is sufficient in revealing the true causal influence or rather just identifying confounders of the Berkeley graduate admission paradox example. The admission probabilities were modeled using the binomial Bayesian framework, and the posterior is approximated when priors are assigned, as illustrated in the notebook in supplementary materials.
@@ 2 Graphs and Properties
Maximal ancestral graphs (MAGs), first introduced by [33], provide a framework for modeling distributions through conditional independence (CI) relations. Compared with directed acyclic graphs (DAGs), MAGs release the assumption of no latent confounders, accommodating data that arises from distributions with more complex independent structures and revealing hidden states in the graphs. These graphs have demonstrated their utility in a variety of contexts, such as inferring causal effects from observational data. While DAGs allow efficient computation of maximum likelihood estimates (MLEs) and scoring (e.g., via BIC), these properties are challenging to extend to ADMGs due to their structural and computational complexity [21].
@@
One can illustrate any causal relationship as an acyclic directed mixed graph (ADMG), which is acyclic and contains only two types of edges: directed ($\rightarrow$) and bidirected ($\leftrightarrow$). A graph $\mathcal{G}$, which consists of a vertex set $\mathcal{V}$ and an edge set $\mathcal{E}$ of pairs of distinct vertices, is acyclic if there does not exist any sequence of vertices $v_{0},v_{1},\ldots,v_{k}$ (with $k\geq 1$) such that:
$\forall i\in\{0,\ldots,k-1\},\ (v_{i},v_{i+1})\in E\quad\text{and}\quad(v_{k},v_{0})\in E.$
In other words, an acyclic $\mathcal{G}$ contains no directed cycles.
For an edge in $\mathcal{E}$ connecting vertices $a$ and $b$, we say these two vertices are the endpoints of the edge and the two vertices are adjacent (if there is no edge between $a$ and $b$, they are nonadjacent). For a vertex $v$ in an ADMG $\mathcal{G}$, we define the parents, siblings, ancestors, and district of $v$, respectively as:
$\text{pa}_{G}(v)=\{w:w\to v\text{ in }G\},$
$\text{sib}_{G}(v)=\{w:w\leftrightarrow v\text{ in }G\},$
$\text{an}_{G}(v)=\{w:w\to\cdots\to v\text{ in }G\text{ or }w=v\},$
$\text{dis}_{G}(v)=\{w:w\leftrightarrow\cdots\leftrightarrow v\text{ in }G\text{ or }w=v\}.$
@@
For an ADMG $\mathcal{G}$, given a subset $W\subseteq\mathcal{V}$, the induced subgraph $G_{W}$ is defined as the graph with vertex set $W$ and edges in $\mathcal{G}$ whose endpoints are both in $W$. Also for the district of a vertex $v$ in an induced subgraph $G_{W}$, we may denote it by $\text{dis}_{W}(v)=\text{dis}_{G_{W}}(v)$.
@@ 3 Brief Introduction To Mixed Integer Quadratic Programming
and $m,n,k,r\in\mathbb{N}$. Equation (1) is often called the cost or loss function, (2) represents the quadratic constraints, (3) are the linear constraints, and $F$ is the set that enforces the integrality constraints for the $r$ components of the decision variable $x$.
@@ 4 Formulation of the Mixed Integer Quadratic Program
If non-zero components of the vector $w_{0}$ are correlated with certain components of $X_{i}$, these components are defined as causal $X$-variables for $Y_{i}$. This means:
@@
For the scenario that there are no exogenous variable perturbations, we describe this with a structural equation model (SEM) as follows:
$Y_{i}$ $\leftarrow X_{i}w_{0}+H_{i}\delta+\epsilon_{Y,i},$ (8)
@@
where $\epsilon_{X,i}$, $\epsilon_{Y,i}$ and $H_{i}$ are mutually independent. $X_{i}$ is the observed $p\times 1$ covariate vector. $H_{i}$ is the unobserved $q\times 1$ confounding variable vector. Our goal is to infer the confounding-free regression parameter $w_{0}$ and stabilize the prediction of the relationship between $Y$ and $X$.
@@ 4.1 Connecting to Causality
The causal parameter $w_{0}$ can be seen as minimizing the worst-case risk:
$w_{0}=\arg\min_{w}\max_{P\in\mathcal{P}}\mathbb{E}_{P}\left[(Y_{i}-X_{i}w)^{2}\right],$
@@
In this perspective, we present the formulation of the Mixed Integer Quadratic Program (MIQP) used to infer the causal structure, with a new binary matrix $B=[b_{j,k}]\in\{0,1\}^{d\times d}$ introduced to account for relationships explained by confounding factors, alongside the weight matrix $W$ and binary adjacency matrix $E$ adopted from the ExDAG model [38]. Weight matrix $W$ represents the model weights. Whenever a field in $W_{j,k}$ is non-zero, the respective value in either $E_{j,k}$ (directed edge) or $B_{j,k}$ is nonzero (bidirected edge). At the same time, we extend the existing formulation by introducing an additional binary input parameter $f_{j,k}$ for each pair of variables $(j,k)$, where $j\neq k$, indicating that there is no direct causal relationship between variables $j$ and $k$, but $j$ and $k$ might have a common cofactor. This follows from the meaning of the edges in a MAG - $\rightarrow$ edge implies a direct causal relationship but does not rule out a possible latent confounding, $\leftrightarrow$ means no direct causal relationship. Formally, we define the Directed Edge Matrix $E$ as
@@
Based on the above assumptions, one can substitute algebraic equalities 9 into 8, and the work of [38] is extended as explained in the next section.

### v2

@@ Abstract
As one transitions from statistical to causal learning, one is seeking the most appropriate causal model. Dynamic Bayesian networks are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by its vertices, and weighted oriented edges suggest the strength of the causal relationships. When there are confounders, one would like to utilize both oriented edges (when the direction of causality is clear) and edges that are not oriented (when there is a confounder or not a relationship), yielding mixed graphs. A little-studied extension of acyclicity to this mixed-graph setting is known as maximally ancestral graphs with consideration of confounders.
@@ 1 Introduction
As one transitions from statistical to causal learning [47], one is seeking the most appropriate causal model. Dynamic Bayesian networks (DBN) [17, 35] are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by its vertices, and weighted oriented edges suggest the strength of the causal relationships. The key challenge in learning DBNs is confounding.
To illustrate the challenge of confounding, let us consider Simpson’s paradox. Simpson’s paradox shows that without considering confounding factors in statistical analysis [34], the direction of causality can be mis-estimated completely. An textbook example [34] comes from the Berkeley graduate admissions [3]. The data show that women find it harder to get admitted to Berkeley graduate schools. Nevertheless, this is because women tend to apply to departments that have lower admission rates. In this example, the choice of the graduate school is the confounder, impacting the probability of admission. Confounding is prevalent throughout high-dimensional statistics [32, 21, e.g.], such as in biomedical sciences.
@@
Specifically, in biomedical sciences, confounders such as socio-economic status, age, or lifestyle factors can distort the true causal relationship between treatments and outcomes [52]. Techniques such as instrumental variables [39, 29], propensity score matching [43], and double machine learning [11] have been widely used to mitigate the effects of confounding in clinical trials and observational studies. In bioinformatics, particularly in genome-wide association studies (GWAS), confounders, including population stratification and environmental exposures, must be controlled to avoid biased estimates in genetic association studies [9, 22]. To mitigate such biases, statistical models that explicitly account for hidden confounders, such as spectral methods and latent variable models, are often employed [22]. Furthermore, meta-analysis and sensitivity analysis are often employed to evaluate the robustness of findings in the presence of potential confounders, especially when combining results from multiple studies [6, 33]. These methodologies ensure that the conclusions drawn are reliable and actionable, improving the credibility of statistical models across disciplines.
@@
In statistical theory, [8] study confounding in detail, and many subsequent works develop further methods. [31] show that leveraging the dominant eigenstructure of time series may improve performance of estimation. Anchor regression, for instance, bridges the gap between causality and robustness by addressing heterogeneity in data [44]. Other significant contributions include spectral deconfounding, a technique designed to mitigate the effects of hidden confounders in high-dimensional settings [7]. This approach provides a framework for robust predictions in the presence of shifts in data distributions. Similarly, the invariance principle has emerged as a cornerstone of causal inference, linking causal structure to robust statistical models [6]. Furthermore, the concept of doubly robust inference offers an alternative framework for addressing hidden confounding factors, combining model robustness with efficiency in high-dimensional scenarios [22]. Together, these developments represent a significant step forward in understanding and addressing the challenges posed by complex causal systems.
While there is a long history of the study of confounding, as suggested above, the extensions of DBNs to allow for confounding are rather more recent. Instead of estimating a Directed Acyclic Graph (DAG), one could estimate a Maximal Ancestral Graph (MAG), cf. [41]. These allow for both direct and indirect relationships among variables modelled as directed and undirected edges, even in the presence of confounding factors. In particular, MAGs can represent feedback loops and bidirectional relationships that DAG-based models, such as DBNs, cannot. This makes MAGs a more powerful tool for capturing the complex dynamics of real-world causal systems.
There are only a few studies of MAG estimation [10, 38, 12, 24, 27, 26, 15]. [40, 49] are applicable to the both discrete and nonparametric cases, which extend DAG to MAG, ADMG or Bow-Free Acyclic Path Diagrams, see more definition in 2. Factorization in MAG is not directly decomposable into individual variables and their parent sets, as in DAGs, but must instead consider components connected by bidirected paths (termed districts or c-components), cf. [40], although [12] proposed to use Markov equivalence classes (MEC) instead. In 2021, [10] introduced a first mixed-integer programming (MIP) formulation, but the number of variables scales with the number of c-components, i. e., exponentially with the number of vertices in the worst case. Such formulations are also known as extended formulations [13]. The same year, [38] explored a constraint-based approach for MAG discovery, leveraging conditional independence testing. Additionally, [53] addresses exogenous covariates in causal formulation that helps explain the heterogeneity in both sampling and causal mechanisms. The dissertation of [24] presented an extension of the imsets of [48] from directed acyclic graphs (DAGs) to towards MAGs [27], which allows for the use of the methods of Studený, and a score-based heuristic [26]. More recently, the paper by [15] enhanced the scalability of methods of [10] by utilizing linear programming (LP) relaxations instead of solving the MIP.
Our approach proposes the first compact formulation of MAG estimation within Mixed-Integer Nonlinear Programming (MINLP), in contrast with the so-called extended formulations of [10, 15]. While both the extended and compact formulations ensure that confounding factors are properly accounted for and the true underlying data-generating process is better represented by the model, our implementation scales further, from 4-5 stochastic processes in the extended formulation to 25 or more stochastic processes with the proposed compact formulation.
@@ 1.1 Motivating Example
Let us revisit the Berkeley graduate admission paradox example of page 1. As in most paradoxes, there is no violation of logic in Simpson’s paradox, just a violation of of intuition. The poor intuition being violated in this case is that a positive association in the entire population should also hold within each department. Overall, females in these data did have a harder time getting admitted to graduate school. But that arose, because female applicants chose the departments that were the most difficult to gain admission to for anyone, male or female. In this example, gender influences the choice of department, and the department influences the chance of admission. Controlling for department reveals a more plausible direct causal influence of gender. If one represents gender as $G$, department as $D$, acceptance as $A$, and academic ability as $U$, the potential confounder, the DAG form is
[…]
Our method, ExMAG, is able to reveal the confounders in this Berkeley graduate admission example, as illustrated in the notebook in supplementary materials and Figure 1.
Figure 1: Performance of ExDBN of [45] (top) and ExMAG (bottom) on the Berkeley graduate admission example. While the DBN identified by ExDBN suggests a causal relationship between gender and admission, ExMAG correctly identifies the confounding.
@@ 2 Graphs and Properties
Maximal ancestral graphs (MAGs), first introduced by [41], provide a framework for modeling distributions through conditional independence (CI) relations. Compared with directed acyclic graphs (DAGs), MAGs allow for latent confounders, accommodating data that arises from distributions with more complex independence structures and revealing hidden states in the graphs. While DAGs allow for the efficient computation of maximum likelihood estimates (MLEs) and scoring (e.g., via BIC), these properties are challenging to extend to MAG due to their structural and computational complexity [25].
Directed Acyclic Graph
A DAG is a directed graph $G=(V,E)$ such that there are no directed cycles. That is, there is no sequence of distinct vertices $v_{1},v_{2},\dots,v_{k}\in V$ such that $(v_{i},v_{i+1})\in E$ for all $11\leq i\leq k-$ 1 and $(v_{k},v_{1})\in E$.
Bow-Free Acyclic Path Diagrams
Let $G$ be a bow-free ADMG (Acyclic Directed Mixed Graph). The graph $G$ is bow-free if and only if there does not exist a directed and bidirected edge between a pair of adjacent vertices. More formally, for any $v_{i},v_{j}\in V$, there cannot exist both $v_{i}\to v_{j}$ (or $v_{i}\leftarrow v_{j}$) and $v_{i}\leftrightarrow v_{j}$ simultaneously, as described by [50] .
@@
Mixed graphs feature two types of edges: directed ($\rightarrow$) and bidirected ($\leftrightarrow$). Mixed graph $\mathcal{G}$ thus consists of a vertex set $\mathcal{V}$, a set of directed edges $\mathcal{E}$ and undirected edges $\mathcal{U}$, where $\mathcal{E}$ are (ordered) pairs of distinct vertices, while $\mathcal{U}$ are (unordered) 2-element subsets of vertices. For a directed edge in $\mathcal{E}$ connecting vertices $v$ and $w$, we say these two vertices are the endpoints of the edge and the two vertices are adjacent (if there is no edge between $v$ and $w$, they are non-adjacent). For a vertex $v$ in $\mathcal{V}$, we define the parents, spouses, ancestors, and district of $v$, respectively as:
$\text{pa}_{\mathcal{G}}(v)=\{w:w\to v\text{ in }\mathcal{G}\},$
$\text{sp}_{\mathcal{G}}(v)=\{w:w\leftrightarrow v\text{ in }\mathcal{G}\}$
$\text{an}_{\mathcal{G}}(v)=\{w:w\to\cdots\to v\text{ in }\mathcal{G}\text{ or }w=v\},$
$\text{dis}_{\mathcal{G}}(v)=\{w:w\leftrightarrow\cdots\leftrightarrow v\text{ in }\mathcal{G}\text{ or }w=v\}.$
Given a directed mixed graph $\mathcal{G}$, the districts define a set of equivalence classes of nodes in $\mathcal{G}$. The district for node $v$ is defined as the connected component of $v$ in the subgraph of $\mathcal{G}$ induced by all bidirected edges. As in a DAG, a mixed graph $\mathcal{G}$ is acyclic if it contains no directed cycles in $\mathcal{E}_{d}$, i.e., an acyclic directed mixed graph (ADMG) [24].
A directed mixed graph G is called an ancestral ADMG if the following condition holds for all pairs of nodes $v$ and $w$ in $\mathcal{G}$:
$\text{If }v\neq w\text{ and }w\in\text{an}_{\mathcal{G}}(v)\cup\text{sp}_{\mathcal{G}}(v),\text{ then }v\notin\text{an}_{\mathcal{G}}(w),$
which is written as, $\mathcal{G}$ is an ancestral ADMG if it contains no directed cycles ($v\to u\to\ldots\to w\to v$) or almost directed cycles [10, 24].
An almost directed cycle is of the form $v\to u\to\ldots\to w\leftrightarrow v$; in other words, $\{v,w\}\in\mathcal{E}_{w}$ is a bidirected edge, and $v\in\text{an}_{\mathcal{G}}(w)$ [10].
Inducing Paths
An inducing path from a variable $X$ to a variable $Y$ in a directed graph $G=(V,E)$ is a path $P=(X=v_{0},v_{1},\dots,v_{n}=Y)$ such that for all intermediate nodes $v_{i}$ (where $11\leq i\leq n-$ 1), $v_{i}\in Z$, where $Z$ is the conditioning set. If the path is blocked by conditioning on $Z$, then it is an inducing path.
A node j on a non-overlapping path is called a collider if contains a non-overlapping subpath (i; j; k) with two arrowheads into j. In mathematical form, collider is represented as
$\text{collider}_{G}(u)=\left\{u:\exists w,v:\begin{array}[]{c}w\to u\leftarrow v\vee\\ w\leftrightarrow u\leftarrow v\vee\\ w\to u\leftrightarrow v\vee\\ w\leftrightarrow u\leftrightarrow v\end{array}\right\}.$
@@
For an ADMG $\mathcal{G}$, given a subset $W\subseteq\mathcal{V}$, the induced subgraph $\mathcal{G}_{W}$ is defined as the graph with vertex set $W$ and edges in $\mathcal{G}$ whose endpoints are both in $W$. Also for the district of a vertex $v$ in an induced subgraph $\mathcal{G}_{W}$, we may denote it by $\text{dis}_{W}(v)=\text{dis}_{G_{W}}(v)$.
@@ 3 Brief Introduction To Mixed Integer Quadratic Programming
and $m,n,k,r\in\mathbb{N}$. In the mathematical form, (1) is often called the cost or loss function, (2) represents the quadratic constraints, (3) is the linear constraints, and $F$ is the set that enforces the integrality constraints for the $r$ components of the decision variable $x$.
@@ 4 Formulation of the Mixed Integer Quadratic Program
  • $H_{i}$ is the unobserved confounding variable vector.
If non-zero components of the vector $\tilde{w}$ are correlated with certain components of $X_{i}$, these components are defined as causal $X$-variables for $Y_{i}$. This means:
@@
For the scenario that there are no exogenous variable perturbations, we describe this with an additive noise model with hidden states as follows:
$Y_{i}$ $\leftarrow X_{i}\tilde{w}+H_{i}\theta+\epsilon_{Y,i},$ (8)
@@
where $\epsilon_{X,i}$, $\epsilon_{Y,i}$ and $H_{i}$ are mutually independent. $X_{i}$ is the observed $p\times 1$ covariate vector. $H_{i}$ is the unobserved $q\times 1$ confounding variable vector. Our goal is to infer the confounding-free regression parameter $\tilde{w}$ and stabilize the prediction of the relationship between $Y$ and $X$.
@@ 4.1 Connecting to Causality
The causal parameter $\tilde{w}$ can be seen as minimizing the worst-case risk:
$\tilde{w}=\arg\min_{w}\max_{P\in\mathcal{P}}\mathbb{E}_{P}\left[(Y_{i}-X_{i}w)^{2}\right],$
@@
In this perspective, we present the formulation of the Mixed Integer Quadratic Program (MIQP) used to infer the causal structure, with a new binary matrix $B=[b_{j,k}]\in\{0,1\}^{d\times d}$ introduced to account for relationships explained by confounding factors, alongside the weight matrix $W$ and binary adjacency matrix $E$ adopted from the ExDAG model [46]. Weight matrix $W$ represents the model weights. Whenever an entry in $W_{j,k}$ is non-zero, the respective value in either $E_{j,k}$ (directed edge) or $B_{j,k}$ is nonzero (bidirected edge). At the same time, we extend the existing formulation by introducing an additional binary input parameter $f_{j,k}$ for each pair of variables $(j,k)$, where $j\neq k$, indicating that there is no direct causal relationship between variables $j$ and $k$, but $j$ and $k$ might have a common cofactor. This follows from the meaning of the edges in a MAG $\rightarrow$ edge implies a direct causal relationship but does not rule out a possible latent confounding, $\leftrightarrow$ means no direct causal relationship. Formally, we define the Directed Edge Matrix $E$ as
@@
Based on these assumptions, one can substitute algebraic equalities 9 into 8, and the work of [46] is extended as explained in the next section.